OpenStudy (anonymous):

"Medal if answered" The surface area of a cylindrical can of radius r and height h is 2πr2 + 2πrh. If the can is three times as high as the radius of its top, express its surface area S as a function of r.

OpenStudy (unklerhaukus):

\[SA = 2\pi r^2+2\pi rh\] And you are given that \(h=3r\) so sub this into \(SA\); replace \(h\) with \(3r\), and simplify

OpenStudy (perl):



OpenStudy (anonymous):

Im not sure those are the correct answers

OpenStudy (anonymous):

\[SA=2\pi r^2+2\pi rh\] \[h=3r\] \[Therefore, SA = 2 \pi r^2 +2 \pi r(3r)=2 \pi r^2 + 6\pi r^2 = 8 \pi r^2\] \[Hence, r = \sqrt{\frac{ SA }{ 8 \pi }}\]

OpenStudy (anonymous):

thank you very much best mathmatician your a life saver !